# Strong Converse Exponent of Quantum Dichotomies

### v1

@@ IV Proof of the achievability part
$\begin{split}&\mathcal{F}_{1}:=\big{\{}(\tau_{1},\tau_{2})~{}|~{}(\tau_{1},\tau_{2})\in\mathcal{S}_{\rho_{1}}\times\mathcal{S}_{\rho_{2}},r<\frac{D(\tau_{1}\|\sigma_{1})}{D(\tau_{2}\|\sigma_{2})}\big{\}},\\ &\mathcal{F}_{2}:=\big{\{}(\tau_{1},\tau_{2})~{}|~{}(\tau_{1},\tau_{2})\in\mathcal{S}_{\rho_{1}}\times\mathcal{S}_{\rho_{2}},r\geq\frac{D(\tau_{1}\|\sigma_{1})}{D(\tau_{2}\|\sigma_{2})}\big{\}}.\end{split}$
@@
Lemma 10
For $\rho_{1},\sigma_{1}\in\mathcal{S}(\mathcal{H}_{1})$ satisfy ${\operatorname{supp}}(\rho_{1})\subset{\operatorname{supp}}(\sigma_{1})$, $\rho_{2},\sigma_{2}\in\mathcal{S}(\mathcal{H}_{2})$ and $r\geq 0$, we have
@@
Proof. According to the definition of $F_{1}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)$, for any $\delta>0$, there exists $(\tau_{1},\tau_{2})\in\mathcal{S}_{\rho_{1}}\times\mathcal{S}_{\rho_{2}}$ such that
@@
$\begin{split}&\mathcal{E}_{n}(\sigma_{1}^{\otimes n})=\sigma_{2}^{\otimes nr},\\ &\lim_{n\rightarrow\infty}\|\mathcal{E}_{n}(\tau_{1}^{\otimes n})-\tau_{2}^{\otimes nr}\|_{1}=0.\end{split}$ (26)
Now, we let $t=4(rD(\tau_{2}\|\rho_{2})+D(\tau_{1}\|\rho_{1}))$, $\epsilon_{n}=2^{-nt}$, $C=\max\{1-r\log\lambda_{\rm{min}}(\rho_{2}),t+r\log|\mathcal{H}_{2}|\}$, where $\lambda_{\rm{min}}(\rho_{2})$ is the smallest non-zero eigenvalue of $\rho_{2}$ and $\rho_{2}^{n}=(1-\epsilon_{n})\rho_{2}^{\otimes nr}+\epsilon_{n}\pi_{n}$, where $\pi_{n}$ is the maximally mixed state onto the subspace ${\operatorname{supp}}(\rho_{2}^{\otimes nr})^{\perp}$. For sufficiently large $n$, $\lambda_{\rm{min}}(\rho_{2}^{n})\geq 2^{-nC}$ and for such $n$, $F(\rho_{2}^{n},\mathcal{E}_{n}(\rho_{1}^{\otimes n}))$ can be evaluated as follows.
$\begin{split}&-\log F^{2}(\rho_{2}^{n},\mathcal{E}_{n}(\rho_{1}^{\otimes n}))\\ \leq&D(\mathcal{E}_{n}(\tau_{1}^{\otimes n})\|\mathcal{E}_{n}(\rho_{1}^{\otimes n}))+D(\mathcal{E}_{n}(\tau_{1}^{\otimes n})\|\rho_{2}^{n})\\ \leq&nD(\tau_{1}\|\rho_{1})+D(\mathcal{E}_{n}(\tau_{1}^{\otimes n})\|\rho_{2}^{n})\\ =&nD(\tau_{1}\|\rho_{1})-H(\mathcal{E}_{n}(\tau_{1}^{\otimes n}))-\operatorname{Tr}\mathcal{E}_{n}(\tau_{1}^{\otimes n})\log\rho_{2}^{n}+H(\tau_{2}^{\otimes nr})+\operatorname{Tr}\tau_{2}^{\otimes nr}\log\rho_{2}^{n}+D(\tau_{2}^{\otimes nr}\|\rho_{2}^{n})\\ =&nD(\tau_{1}\|\rho_{1})-H(\mathcal{E}_{n}(\tau_{1}^{\otimes n}))+H(\tau_{2}^{\otimes nr})+\operatorname{Tr}\big{(}\tau_{2}^{\otimes nr}-\mathcal{E}_{n}(\tau_{1}^{\otimes n})\big{)}\log\rho_{2}^{n}+D(\tau_{2}^{\otimes nr}\|\rho_{2}^{n})\\ \leq&nD(\tau_{1}\|\rho_{1})+\frac{1}{2}\|\mathcal{E}_{n}(\tau_{1}^{\otimes n})-\tau_{2}^{\otimes nr}\|_{1}\log|\mathcal{H}_{2}|^{nr}+h(\frac{1}{2}\|\mathcal{E}_{n}(\tau_{1}^{\otimes n})-\tau_{2}^{\otimes nr}\|_{1})\\ &+\|\mathcal{E}_{n}(\tau_{1}^{\otimes n})-\tau_{2}^{\otimes nr}\|_{1}\|\log\rho_{2}^{n}\|_{\infty}+D(\tau_{2}^{\otimes nr}\|\rho_{2}^{n})\\ \leq&nD(\tau_{1}\|\rho_{1})+\frac{1}{2}\|\mathcal{E}_{n}(\tau_{1}^{\otimes n})-\tau_{2}^{\otimes nr}\|_{1}\log|\mathcal{H}_{2}|^{nr}+h(\frac{1}{2}\|\mathcal{E}_{n}(\tau_{1}^{\otimes n})-\tau_{2}^{\otimes nr}\|_{1})\\ &+nC\|\mathcal{E}_{n}(\tau_{1}^{\otimes n})-\tau_{2}^{\otimes nr}\|_{1}+D(\tau_{2}^{\otimes nr}\|\rho_{2}^{n})\\ \leq&nD(\tau_{1}\|\rho_{1})+\frac{1}{2}\|\mathcal{E}_{n}(\tau_{1}^{\otimes n})-\tau_{2}^{\otimes nr}\|_{1}\log|\mathcal{H}_{2}|^{nr}+h(\frac{1}{2}\|\mathcal{E}_{n}(\tau_{1}^{\otimes n})-\tau_{2}^{\otimes nr}\|_{1})\\ &+nC\|\mathcal{E}_{n}(\tau_{1}^{\otimes n})-\tau_{2}^{\otimes nr}\|_{1}+nrD(\tau_{2}\|\rho_{2})-\log(1-\epsilon_{n}),\end{split}$ (27)
where the second line comes from Lemma 15 in Appendix, the third line is due to the data processing inequality, the sixth line follows from the Fannes-Audenaert inequality and the operator Hölder inequality. Dividing $n$ in both sides of Eq. (27) and taking the limit of $n$, we have
@@
Eq. (28), Eq. (29) and the definition of $\epsilon_{n}$ gives that
$\lim_{n\rightarrow\infty}-\frac{\log F^{2}(\rho_{2}^{\otimes nr},\mathcal{E}_{n}(\rho_{1}^{\otimes n}))}{n}\leq D(\tau_{1}\|\rho_{1})+rD(\tau_{2}\|\rho_{2})\leq F_{1}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)+\delta.$ (30)
Eq. (30) and Remark 4 implies that
@@
[…]
Lemma 11
For $\rho_{1},\sigma_{1}\in\mathcal{S}(\mathcal{H}_{1})$ satisfy ${\operatorname{supp}}(\rho_{1})\subset{\operatorname{supp}}(\sigma_{1})$, $\rho_{2},\sigma_{2}\in\mathcal{S}(\mathcal{H}_{2})$ and $r\geq 0$, we have
@@
Proof. According to the definition of $F_{2}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)$, there exists $(\tau_{1},\tau_{2})\in\mathcal{S}_{\rho_{1}}\times\mathcal{S}_{\rho_{2}}$ such that
@@
$\begin{split}&\lim_{n\rightarrow\infty}-\frac{\log F^{2}(\rho_{2}^{\otimes nr},\Phi_{n}(\rho_{1}^{\otimes n}))}{n}\\ =&\lim_{n\rightarrow\infty}-\big{\{}\frac{\log F^{2}(\rho_{2}^{\otimes nr_{1}},\mathcal{E}_{n}(\rho_{1}^{\otimes n}))}{n}+\frac{\log F^{2}(\rho_{2}^{\otimes n(r-r_{1})},\sigma_{2}^{\otimes n(r-r_{1})})}{n}\big{\}}\\ \leq&D(\tau_{1}\|\rho_{1})+r_{1}D(\tau_{2}\|\rho_{2})-(r-r_{1})\log F^{2}(\rho_{2},\sigma_{2})\\ \leq&D(\tau_{1}\|\rho_{1})+r_{1}D(\tau_{2}\|\rho_{2})+(r-r_{1})\{D(\tau_{2}\|\rho_{2})+D(\tau_{2}\|\sigma_{2})\}\\ =&D(\tau_{1}\|\rho_{1})+rD(\tau_{2}\|\rho_{2})+rD(\tau_{2}\|\sigma_{2})-D(\tau_{1}\|\sigma_{1})+\delta D(\tau_{2}\|\sigma_{2}),\end{split}$ (37)
@@
[…]
@@
Proof of the achievability part of Eq. (14). If ${\operatorname{supp}}(\rho_{1})\subseteq{\operatorname{supp}}(\sigma_{1})$ does not hold, it is obvious that all terms in Eq. (14) are equal to 00 and the assertion holds trivially. Hence, for the rest we assume that ${\operatorname{supp}}(\rho_{1})\subseteq{\operatorname{supp}}(\sigma_{1})$. We first fix an integer $m$ and let $\rho_{1}^{m}=\mathcal{E}_{\sigma_{1}^{\otimes m}}(\rho_{1}^{\otimes m})$, $\rho_{2}^{m}=\mathcal{E}_{\sigma_{2}^{\otimes m}}(\rho_{2}^{\otimes m})$. From Lemma 9 , we know that there exists a sequence of quantum channels $\{\Phi_{k}\}_{k\in\mathbb{N}}$ such that
@@
$\begin{split}&\lim_{k\rightarrow\infty}-\frac{\log F^{2}({\rho_{2}}^{\otimes mkr},\mathcal{E}_{\sigma_{2}^{\otimes m}}^{\otimes kr}\Phi_{k}({\rho^{m}_{1}}^{\otimes k}))}{k}\\ \leq&\sup_{\frac{1}{2}<\alpha<1}\frac{1-\alpha}{\alpha}\big{\{}rD_{\alpha}^{\flat}(\rho^{m}_{2}\|\sigma^{\otimes m}_{2})-D_{\beta}^{\flat}(\rho^{m}_{1}\|\sigma^{\otimes m}_{1})\big{\}}+r\log v(\sigma_{2}^{\otimes m}).\end{split}$ (41)
@@
$\mathcal{E}_{n}(X)=\mathcal{E}_{\sigma_{2}^{\otimes m}}^{\otimes kr}\Phi_{k}\mathcal{E}_{\sigma_{1}^{\otimes m}}^{\otimes k}\operatorname{Tr}_{l}(X)\otimes\sigma_{2}^{\otimes lr},$
@@
$\begin{split}&\lim_{n\rightarrow\infty}-\frac{\log F^{2}({\rho_{2}}^{\otimes nr},\mathcal{E}_{n}(\rho_{1}^{\otimes n}))}{n}\\ =&\lim_{k\rightarrow\infty}-\frac{\log F^{2}({\rho_{2}}^{\otimes mkr},\mathcal{E}_{\sigma_{2}^{\otimes m}}^{\otimes kr}\Phi_{k}({\rho^{m}_{1}}^{\otimes k}))F^{2}(\rho_{2}^{\otimes lr},\sigma_{2}^{\otimes lr})}{mk+l}\\ \leq&\sup_{\frac{1}{2}<\alpha<1}\frac{1-\alpha}{\alpha}\big{\{}r\frac{D_{\alpha}^{\flat}(\rho^{m}_{2}\|\sigma^{\otimes m}_{2})}{m}-\frac{D_{\beta}^{\flat}(\rho^{m}_{1}\|\sigma^{\otimes m}_{1})}{m}\big{\}}+r\frac{\log v(\sigma_{2}^{\otimes m})}{m}.\end{split}$ (43)
@@
$\begin{split}&E^{P}_{sc}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)\\ \leq&\sup_{\frac{1}{2}<\alpha<1}\frac{1-\alpha}{\alpha}\big{\{}r\frac{D_{\alpha}^{\flat}(\rho^{m}_{2}\|\sigma^{\otimes m}_{2})}{m}-\frac{D_{\beta}^{\flat}(\rho^{m}_{1}\|\sigma^{\otimes m}_{1})}{m}\big{\}}+r\frac{\log v(\sigma_{2}^{\otimes m})}{m}\\ =&\sup_{\frac{1}{2}<\alpha<1}\frac{1-\alpha}{\alpha}\big{\{}r\frac{D_{\alpha}^{*}(\rho^{m}_{2}\|\sigma^{\otimes m}_{2})}{m}-\frac{D_{\beta}^{*}(\rho^{m}_{1}\|\sigma^{\otimes m}_{1})}{m}\big{\}}+r\frac{\log v(\sigma_{2}^{\otimes m})}{m}\\ \leq&\sup_{\frac{1}{2}<\alpha<1}\frac{1-\alpha}{\alpha}\big{\{}rD_{\alpha}^{*}(\rho_{2}\|\sigma_{2})-D_{\beta}^{*}(\rho_{1}\|\sigma_{1})+2\frac{v(\sigma_{1}^{\otimes m})}{m}\big{\}}+r\frac{\log v(\sigma_{2}^{\otimes m})}{m}\\ \leq&\sup_{\frac{1}{2}<\alpha<1}\frac{1-\alpha}{\alpha}\big{\{}rD_{\alpha}^{*}(\rho_{2}\|\sigma_{2})-D_{\beta}^{*}(\rho_{1}\|\sigma_{1})\big{\}}+2\frac{v(\sigma_{1}^{\otimes m})}{m}+r\frac{\log v(\sigma_{2}^{\otimes m})}{m}\\ =&\sup_{\frac{1}{2}\leq\alpha\leq 1}\frac{1-\alpha}{\alpha}\big{\{}rD_{\alpha}^{*}(\rho_{2}\|\sigma_{2})-D_{\frac{\alpha}{2\alpha-1}}^{*}(\rho_{1}\|\sigma_{1})\big{\}}+2\frac{v(\sigma_{1}^{\otimes m})}{m}+r\frac{\log v(\sigma_{2}^{\otimes m})}{m},\end{split}$ (44)
[…]
V Proof of the optimality part
In this section, we prove the optimality part of Eq. (14). The proof is essentially from a easy application of the operator Hölder inequality.
Proof of the optimality part of Eq. (14). If ${\operatorname{supp}}(\rho_{1})\subseteq{\operatorname{supp}}(\sigma_{1})$ does not hold, it is obvious that all terms in Eq. (14) are equal to 00 and the assertion holds trivially. Hence, for the rest we assume that ${\operatorname{supp}}(\rho_{1})\subseteq{\operatorname{supp}}(\sigma_{1})$. By Lemma 16 in Appendix, for any quantum channel $\mathcal{E}_{n}$ such that $\mathcal{E}_{n}(\sigma_{1}^{\otimes n})=\sigma_{2}^{\otimes nr}$ and any $\alpha\in(\frac{1}{2},1)$, $\beta=\frac{\alpha}{2\alpha-1}$, we have
$\begin{split}&\frac{2\alpha}{1-\alpha}\log F\big{(}\mathcal{E}_{n}(\rho_{1}^{\otimes n}),\rho_{2}^{\otimes nr}\big{)}\\ \leq&-D_{\alpha}^{*}(\rho_{2}^{\otimes nr}\|\sigma_{2}^{\otimes nr})+D^{*}_{\beta}(\mathcal{E}_{n}(\rho_{1}^{\otimes n})\|\sigma_{2}^{\otimes nr})\\ =&-D_{\alpha}^{*}(\rho_{2}^{\otimes nr}\|\sigma_{2}^{\otimes nr})+D^{*}_{\beta}(\mathcal{E}_{n}(\rho_{1}^{\otimes n})\|\mathcal{E}_{n}(\sigma_{1}^{\otimes n}))\\ \leq&-nrD_{\alpha}^{*}(\rho_{2}\|\sigma_{2})+nD_{\beta}^{*}(\rho_{1}\|\sigma_{1}).\end{split}$ (45)
Eq. (45) implies that
$\begin{split}E^{P}_{sc}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)&=\lim_{n\rightarrow\infty}\frac{-1}{n}\log\max_{\mathcal{E}_{n}:\mathcal{E}_{n}(\sigma_{1}^{\otimes n})=\sigma_{2}^{\otimes nr}}F^{2}\big{(}\mathcal{E}_{n}(\rho_{1}^{\otimes n}),\rho_{2}^{\otimes nr}\big{)}\\ &\geq\frac{1-\alpha}{\alpha}\big{\{}rD_{\alpha}^{*}(\rho_{2}\|\sigma_{2})-D_{\beta}^{*}(\rho_{1}\|\sigma_{1})\big{\}}.\end{split}$ (46)
Noticing that Eq. (46) holds for any $\alpha\in(\frac{1}{2},1)$, we have
$\begin{split}E^{P}_{sc}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)&\geq\sup_{\frac{1}{2}<\alpha<1}\frac{1-\alpha}{\alpha}\big{\{}rD_{\alpha}^{*}(\rho_{2}\|\sigma_{2})-D_{\beta}^{*}(\rho_{1}\|\sigma_{1})\big{\}}\\ &=\sup_{\frac{1}{2}\leq\alpha\leq 1}\frac{1-\alpha}{\alpha}\big{\{}rD_{\alpha}^{*}(\rho_{2}\|\sigma_{2})-D_{\beta}^{*}(\rho_{1}\|\sigma_{1})\big{\}}\\ &=\sup_{\frac{1}{2}\leq\alpha\leq 1}\frac{1-\alpha}{\alpha}\big{\{}rD_{\alpha}^{*}(\rho_{2}\|\sigma_{2})-D_{\frac{\alpha}{2\alpha-1}}^{*}(\rho_{1}\|\sigma_{1})\big{\}}.\end{split}$ (47)
[…]
VI Discussion

### v2

@@ IV Proof of the achievability part
$\begin{split}&\mathcal{F}_{1}:=\left\{(\tau_{1},\tau_{2})~{}|~{}(\tau_{1},\tau_{2})\in\mathcal{S}_{\rho_{1}}\times\mathcal{S}_{\rho_{2}},r<\frac{D(\tau_{1}\|\sigma_{1})}{D(\tau_{2}\|\sigma_{2})}\right\},\\ &\mathcal{F}_{2}:=\left\{(\tau_{1},\tau_{2})~{}|~{}(\tau_{1},\tau_{2})\in\mathcal{S}_{\rho_{1}}\times\mathcal{S}_{\rho_{2}},r\geq\frac{D(\tau_{1}\|\sigma_{1})}{D(\tau_{2}\|\sigma_{2})}\right\}.\end{split}$
@@
Lemma 8.
For $\rho_{1},\sigma_{1}\in\mathcal{S}(\mathcal{H}_{1})$ that satisfy $\mathop{\rm supp}(\rho_{1})\subset\mathop{\rm supp}(\sigma_{1})$, $\rho_{2},\sigma_{2}\in\mathcal{S}(\mathcal{H}_{2})$ and $r\geq 0$, we have
@@
Proof: According to the definition of $F_{1}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)$, for any $\delta>0$, there exists $(\tau_{1},\tau_{2})\in\mathcal{S}_{\rho_{1}}\times\mathcal{S}_{\rho_{2}}$ such that
@@
$\begin{split}&\mathcal{E}_{n}(\sigma_{1}^{{\otimes}n})=\sigma_{2}^{{\otimes}nr},\\ \lim_{n\rightarrow\infty}&\|\mathcal{E}_{n}(\tau_{1}^{{\otimes}n})-\tau_{2}^{{\otimes}nr}\|_{1}=0.\end{split}$ (27)
Now, we let $t=4(rD(\tau_{2}\|\rho_{2})+D(\tau_{1}\|\rho_{1}))$, $\epsilon_{n}=2^{-nt}$, $C=\max\{1-r\log\lambda_{\rm{min}}(\rho_{2}),t+r\log|\mathcal{H}_{2}|\}$, where $\lambda_{\rm{min}}(\rho_{2})$ is the smallest non-zero eigenvalue of $\rho_{2}$ and $\rho_{2}^{n}=(1-\epsilon_{n})\rho_{2}^{{\otimes}nr}+\epsilon_{n}\pi_{n}$, where $\pi_{n}$ is the maximally mixed state onto the subspace $\mathop{\rm supp}(\rho_{2}^{{\otimes}nr})^{\perp}$. For sufficiently large $n$, $\lambda_{\rm{min}}(\rho_{2}^{n})\geq 2^{-nC}$ and for such $n$, $F(\rho_{2}^{n},\mathcal{E}_{n}(\rho_{1}^{{\otimes}n}))$ can be evaluated as follows.
$\begin{split}&-\log F^{2}(\rho_{2}^{n},\mathcal{E}_{n}(\rho_{1}^{{\otimes}n}))\\ \leq&D(\mathcal{E}_{n}(\tau_{1}^{{\otimes}n})\|\mathcal{E}_{n}(\rho_{1}^{{\otimes}n}))+D(\mathcal{E}_{n}(\tau_{1}^{{\otimes}n})\|\rho_{2}^{n})\\ \leq&nD(\tau_{1}\|\rho_{1})+D(\mathcal{E}_{n}(\tau_{1}^{{\otimes}n})\|\rho_{2}^{n})\\ =&nD(\tau_{1}\|\rho_{1})-H(\mathcal{E}_{n}(\tau_{1}^{{\otimes}n}))-\operatorname{Tr}\mathcal{E}_{n}(\tau_{1}^{{\otimes}n})\log\rho_{2}^{n}+H(\tau_{2}^{{\otimes}nr})+\operatorname{Tr}\tau_{2}^{{\otimes}nr}\log\rho_{2}^{n}+D(\tau_{2}^{{\otimes}nr}\|\rho_{2}^{n})\\ =&nD(\tau_{1}\|\rho_{1})-H(\mathcal{E}_{n}(\tau_{1}^{{\otimes}n}))+H(\tau_{2}^{{\otimes}nr})+\operatorname{Tr}\big{(}\tau_{2}^{{\otimes}nr}-\mathcal{E}_{n}(\tau_{1}^{{\otimes}n})\big{)}\log\rho_{2}^{n}+D(\tau_{2}^{{\otimes}nr}\|\rho_{2}^{n}),\end{split}$ (28)
where the second line comes from Lemma 13 in Appendix, the third line is due to the data processing inequality. By making use of Fannes-Audenaert inequality and operator Hölder inequality, we can further upper bound Eq. (28) as
$\begin{split}&-\log F^{2}(\rho_{2}^{n},\mathcal{E}_{n}(\rho_{1}^{{\otimes}n}))\\ \leq&nD(\tau_{1}\|\rho_{1})+\frac{1}{2}\|\mathcal{E}_{n}(\tau_{1}^{{\otimes}n})-\tau_{2}^{{\otimes}nr}\|_{1}\log|\mathcal{H}_{2}|^{nr}+h(\frac{1}{2}\|\mathcal{E}_{n}(\tau_{1}^{{\otimes}n})-\tau_{2}^{{\otimes}nr}\|_{1})\\ &+\|\mathcal{E}_{n}(\tau_{1}^{{\otimes}n})-\tau_{2}^{{\otimes}nr}\|_{1}\|\log\rho_{2}^{n}\|_{\infty}+D(\tau_{2}^{{\otimes}nr}\|\rho_{2}^{n})\\ \leq&nD(\tau_{1}\|\rho_{1})+\frac{1}{2}\|\mathcal{E}_{n}(\tau_{1}^{{\otimes}n})-\tau_{2}^{{\otimes}nr}\|_{1}\log|\mathcal{H}_{2}|^{nr}+h(\frac{1}{2}\|\mathcal{E}_{n}(\tau_{1}^{{\otimes}n})-\tau_{2}^{{\otimes}nr}\|_{1})\\ &+nC\|\mathcal{E}_{n}(\tau_{1}^{{\otimes}n})-\tau_{2}^{{\otimes}nr}\|_{1}+nrD(\tau_{2}\|\rho_{2})-\log(1-\epsilon_{n}).\end{split}$ (29)
Dividing $n$ in both sides of Eq. (29) and taking the limit of $n$, we have
@@
Eq. (30), Eq. (31) and the definition of $\epsilon_{n}$ give that
$\lim_{n\rightarrow\infty}-\frac{\log F^{2}(\mathcal{E}_{n}(\rho_{1}^{{\otimes}n}),\rho_{2}^{{\otimes}nr})}{n}\leq D(\tau_{1}\|\rho_{1})+rD(\tau_{2}\|\rho_{2})\leq F_{1}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)+\delta.$ (32)
Eq. (32) and Remark 1 imply that
@@
Because Eq. (33) holds for any $\delta>0$, by letting $\delta\rightarrow 0$, we complete the proof.
Lemma 9.
For $\rho_{1},\sigma_{1}\in\mathcal{S}(\mathcal{H}_{1})$ that satisfy $\mathop{\rm supp}(\rho_{1})\subset\mathop{\rm supp}(\sigma_{1})$, $\rho_{2},\sigma_{2}\in\mathcal{S}(\mathcal{H}_{2})$ and $r\geq 0$, we have
@@
Proof: According to the definition of $F_{2}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)$, there exists $(\tau_{1},\tau_{2})\in\mathcal{S}_{\rho_{1}}\times\mathcal{S}_{\rho_{2}}$ such that
@@
$\begin{split}&\lim_{n\rightarrow\infty}-\frac{\log F^{2}(\rho_{2}^{{\otimes}nr},\Phi_{n}(\rho_{1}^{{\otimes}n}))}{n}\\ =&\lim_{n\rightarrow\infty}-\bigg{\{}\frac{\log F^{2}(\rho_{2}^{{\otimes}nr_{1}},\mathcal{E}_{n}(\rho_{1}^{{\otimes}n}))}{n}+\frac{\log F^{2}(\rho_{2}^{{\otimes}n(r-r_{1})},\sigma_{2}^{{\otimes}n(r-r_{1})})}{n}\bigg{\}}\\ \leq&D(\tau_{1}\|\rho_{1})+r_{1}D(\tau_{2}\|\rho_{2})-(r-r_{1})\log F^{2}(\rho_{2},\sigma_{2})\\ \leq&D(\tau_{1}\|\rho_{1})+r_{1}D(\tau_{2}\|\rho_{2})+(r-r_{1})\left\{D(\tau_{2}\|\rho_{2})+D(\tau_{2}\|\sigma_{2})\right\}\\ =&D(\tau_{1}\|\rho_{1})+rD(\tau_{2}\|\rho_{2})+rD(\tau_{2}\|\sigma_{2})-D(\tau_{1}\|\sigma_{1})+\delta D(\tau_{2}\|\sigma_{2}),\end{split}$ (39)
@@
Noticing that Eq. (40) holds for any $\delta>0$, by letting $\delta\rightarrow 0$, we complete the proof.
@@
Proof of the achievability part of Eq. (15):   If $\mathop{\rm supp}(\rho_{1})\subseteq\mathop{\rm supp}(\sigma_{1})$ does not hold, it is obvious that all terms in Eq. (15) equal to 00 and the assertion holds trivially. Hence, for the rest we assume that $\mathop{\rm supp}(\rho_{1})\subseteq\mathop{\rm supp}(\sigma_{1})$. We first fix an integer $m$ and let $\rho_{1}^{m}=\mathcal{E}_{\sigma_{1}^{{\otimes}m}}(\rho_{1}^{{\otimes}m})$, $\rho_{2}^{m}=\mathcal{E}_{\sigma_{2}^{{\otimes}m}}(\rho_{2}^{{\otimes}m})$. From Lemma 7 , we know that there exists a sequence of quantum channels $\{\Phi_{k}\}_{k\in\mathbb{N}}$ such that
@@
$\begin{split}&\lim_{k\rightarrow\infty}-\frac{\log F^{2}({\rho_{2}}^{{\otimes}mkr},\mathcal{E}_{\sigma_{2}^{{\otimes}m}}^{{\otimes}kr}\circ\Phi_{k}({\rho^{m}_{1}}^{{\otimes}k}))}{k}\\ \leq&\lim_{k\rightarrow\infty}-\frac{\log v^{-kr}(\sigma_{2}^{{\otimes}m})F^{2}({\rho^{m}_{2}}^{{\otimes}kr},\mathcal{E}_{\sigma_{2}^{{\otimes}m}}^{{\otimes}kr}\circ\Phi_{k}({\rho^{m}_{1}}^{{\otimes}k}))}{k}\\ \leq&\lim_{k\rightarrow\infty}-\frac{\log v^{-kr}(\sigma_{2}^{{\otimes}m})F^{2}({\rho^{m}_{2}}^{{\otimes}kr},\Phi_{k}({\rho^{m}_{1}}^{{\otimes}k}))}{k}\\ \leq&\sup_{\frac{1}{2}<\alpha<1}\frac{1-\alpha}{\alpha}\big{\{}rD_{\alpha}^{\flat}(\rho^{m}_{2}\|\sigma^{{\otimes}m}_{2})-D_{\beta}^{\flat}(\rho^{m}_{1}\|\sigma^{{\otimes}m}_{1})\big{\}}+r\log v(\sigma_{2}^{{\otimes}m}).\end{split}$ (43)
@@
$\mathcal{E}_{n}(X)=\big{(}\mathcal{E}_{\sigma_{2}^{{\otimes}m}}^{{\otimes}kr}\circ\Phi_{k}\circ\mathcal{E}_{\sigma_{1}^{{\otimes}m}}^{{\otimes}k}\circ\operatorname{Tr}_{l}(X)\big{)}{\otimes}\sigma_{2}^{{\otimes}lr},$
@@
$\begin{split}&\lim_{n\rightarrow\infty}-\frac{\log F^{2}({\rho_{2}}^{{\otimes}nr},\mathcal{E}_{n}(\rho_{1}^{{\otimes}n}))}{n}\\ =&\lim_{k\rightarrow\infty}-\frac{\log F^{2}({\rho_{2}}^{{\otimes}mkr},\mathcal{E}_{\sigma_{2}^{{\otimes}m}}^{{\otimes}kr}\circ\Phi_{k}({\rho^{m}_{1}}^{{\otimes}k}))F^{2}(\rho_{2}^{{\otimes}lr},\sigma_{2}^{{\otimes}lr})}{mk+l}\\ \leq&\sup_{\frac{1}{2}<\alpha<1}\frac{1-\alpha}{\alpha}\bigg{\{}r\frac{D_{\alpha}^{\flat}(\rho^{m}_{2}\|\sigma^{{\otimes}m}_{2})}{m}-\frac{D_{\beta}^{\flat}(\rho^{m}_{1}\|\sigma^{{\otimes}m}_{1})}{m}\bigg{\}}+r\frac{\log v(\sigma_{2}^{{\otimes}m})}{m}.\end{split}$ (45)
@@
$\begin{split}&E^{P}_{sc}(\rho_{1},\sigma_{1},\rho_{2},\sigma_{2},r)\\ \leq&\sup_{\frac{1}{2}<\alpha<1}\frac{1-\alpha}{\alpha}\left\{r\frac{D_{\alpha}^{\flat}(\rho^{m}_{2}\|\sigma^{{\otimes}m}_{2})}{m}-\frac{D_{\beta}^{\flat}(\rho^{m}_{1}\|\sigma^{{\otimes}m}_{1})}{m}\right\}+r\frac{\log v(\sigma_{2}^{{\otimes}m})}{m}\\ =&\sup_{\frac{1}{2}<\alpha<1}\frac{1-\alpha}{\alpha}\left\{r\frac{D_{\alpha}^{*}(\rho^{m}_{2}\|\sigma^{{\otimes}m}_{2})}{m}-\frac{D_{\beta}^{*}(\rho^{m}_{1}\|\sigma^{{\otimes}m}_{1})}{m}\right\}+r\frac{\log v(\sigma_{2}^{{\otimes}m})}{m}\\ \leq&\sup_{\frac{1}{2}<\alpha<1}\frac{1-\alpha}{\alpha}\left\{rD_{\alpha}^{*}(\rho_{2}\|\sigma_{2})-D_{\beta}^{*}(\rho_{1}\|\sigma_{1})+2\frac{v(\sigma_{1}^{{\otimes}m})}{m}\right\}+r\frac{\log v(\sigma_{2}^{{\otimes}m})}{m}\\ \leq&\sup_{\frac{1}{2}\leq\alpha\leq 1}\frac{1-\alpha}{\alpha}\left\{rD_{\alpha}^{*}(\rho_{2}\|\sigma_{2})-D_{\beta}^{*}(\rho_{1}\|\sigma_{1})\right\}+2\frac{v(\sigma_{1}^{{\otimes}m})}{m}+r\frac{\log v(\sigma_{2}^{{\otimes}m})}{m}\\ =&\sup_{\frac{1}{2}\leq\alpha\leq 1}\frac{1-\alpha}{\alpha}\big{\{}rD_{\alpha}^{*}(\rho_{2}\|\sigma_{2})-D_{\frac{\alpha}{2\alpha-1}}^{*}(\rho_{1}\|\sigma_{1})\big{\}}+2\frac{v(\sigma_{1}^{{\otimes}m})}{m}+r\frac{\log v(\sigma_{2}^{{\otimes}m})}{m},\end{split}$ (46)
where the third line is because that $\rho^{m}_{1}$ and $\rho^{m}_{2}$ commute with $\sigma_{1}^{{\otimes}m}$ and $\sigma_{2}^{{\otimes}m}$, respectively, the fourth line is due to the data processing inequality and Proposition 2 (v). Because Eq. (46) holds for any $m$, by letting $m\rightarrow\infty$, we complete the proof. ∎
V Discussion